CONFORMATIONAL ISOMERIZATION OF bis-(4-HYDROXYPHENYL)METHANE IN A SUPERSONIC JET EXPANSION. PART II: INTERNAL MIXING AND LOW BARRIER POTENTIAL ENERGY SURFACE IN THE S_{1} STATE.

\underline{CHRISTIAN W. MÜLLER}, CHIRANTHA P. RODRIGO, WILLIAM H. JAMES III, NATHAN R. PILLSBURY and TIMOTHY S. ZWIER, Department of Chemistry, Purdue University, West Lafayette, IN 47907-2084.

The close proximity of two identical ultraviolet chromophores render bis-(4-hydroxyphenyl)methane (B4HPM) an interesting case for the study of the dependence of excitonic coupling on the distortion along low-frequency largeamplitude vibrational coordinates, in particular the phenol ring torsional coordinates present in B4HPM. We have studied the fluorescence excitation spectrum, the UV-UV holeburning spectra and several single vibronic level fluorescence spectra of the Ã^{1}B(S_{1}) ← X̃^{1}A(S_{0}) and the B̃^{1}A(S_{2}) ← X̃^{1}A(S_{0}) transition of all three conformers of B4HPM in a supersonic jet. Excitonic splitting between the two chromophores shifts the second excited state S_{2} ← S_{0} by merely 132 cm^{−1} from the S_{1} ← S_{0} origin in both symmetric conformers. The analysis of the dispersed fluorescence spectra of the S_{2} origins reveals that these levels are internally mixed with nearby S_{1} vibronic levels, providing a fingerprint of the levels involved in the mixing. The dispersed fluorescence spectra of several low-energy S_{1} ← S_{0} vibronic transitions of a specific conformer were taken under systematic variation of the collision frequency in the region where supersonic jet and laser pulse train intersect. These spectra reveal fluorescence contributions from the other two conformers, thus indicating the presence of low-energy conformational barriers (∼ 40–80 cm^{−1}) in the S_{1} state of B4HPM.